# On irrationals with Lagrange value exactly 33

### v2

@@ 1. Introduction
1.1. Lagrange Spectrum and bi-infinite sequences
@@ 1.2. Statement of the main result
A classical result by Markov Markoff1879 ,Markoff1880 from 1880’s states that $\mathcal{L}$ and $\mathcal{M}$ coincide before $3$ and this initial part is exactly a discrete sequence accumulating at $3$, which is
@@
where $l_{1}<l_{2}<l_{3}<\dots$ is an increasing sequence of positive integers and we can show that those $y$ have Lagrange value $3$.
Remark 1.1.
@@
It is relatively easy to give concrete examples of sequences in $m^{-1}(3)$, but these “easy” examples are periodic at both sides and have Lagrange value less than $3$. Sequences in $m^{-1}(3)\cap l^{-1}(3)$ are not so direct to describe, nevertheless with the help of Bombieri’s characterization we have a criteria (see Lemma 2.22) to produce quite a lot such elements.
@@
A detailed sketch of proof is provided at the beginning of Section $3$. We first construct uncountably many elements in the set $Y_{3}(\infty)$ and $Y_{3}(0)$. Then by gluing some carefully chosen finite words at the beginning of elements in $Y_{3}(0)$, we can map $Y_{3}(0)$ injectively into $Y_{3}(n)$ for $0<n<\infty$.
@@ 2.1. Markov’s Theorem
A well known uniqueness conjecture by Frobenius states that all the Markov numbers have multiplicity of $1$ and the conjecture still remains open.
Markov’s theorem gives a complete description of the spectrum $\mathcal{L}$ and $\mathcal{M}$ at the part before $3$. Here we state a version in Bom .
@@
2.2. Structure of Admissible Bi-infinite Words
@@
If there is one letter bigger than $2$ appearing in a bi-infinite word, then the Markov value will be strictly bigger than $3$, so the only possible letters that could appear in an admissible bi-infinite word are $1$ and $2$. Moreover, it was shown in (Bom, , Lemma 9) that the number of consecutive $1$s or $2$s in an admissible bi-infinite word is always even or infinite, so an admissible bi-infinite word can be written as a bi-infinite word of alphabet $\{a=22,b=11\}$. In (Bom, , Lemma 11), admissible bi-infinite words were classified into four possible types:
@@
We only consider positive words, that is, words with positive exponents. If the word is no longer positive after applying $u$ or $v$, then we say that the substitution $u$ or $v$ is not well defined. A remark is that if $u$ or $v$ is well defined for some bi-infinite word $\underline{\omega}$, then all the exponents of $a$ or $b$ in the word $\underline{\omega}$ should be $1$.
@@
We will focus on cuts with value bigger than $3$ or not bigger than $3$.
@@
A cut of an infinite or bi-infinite word is bad if the cut has value strictly bigger than $3$, otherwise the cut is good.
@@
For every $n\in\mathbb{N}\cup\{\infty\}$, the elements in $X_{3}(n)$ are infinite words with Lagrange value $3$ and with exactly $n$ bad cuts. As was stated in Section 1, our main idea is to take projections of elements in $m^{-1}(3)\cap l^{-1}(3)$, and we have to control the number of bad cuts inside the projections.
@@
An important observation is that good cuts for a bi-infinite word $\underline{\omega}$ can change to bad for the infinite word $\pi(\underline{\omega})$. For example, the bi-infinite word $(abb)^{\infty}$ has only good cuts (since it has Markov value less than $3$) but the projection $abbabb\dots$ has bad cuts, because
@@ 2.3. Renormalization
A non-constant bi-infinite word $\underline{\omega}$ is strongly admissible if and only if the renormalization algorithm stops at some finite step and reaches $\underline{\omega}_{0}$, and a non-constant bi-infinite word has Markov value exactly $3$ if and only if the renormalization algorithm never stops.
@@
The Cohn tree $T$ is an infinite binary tree generated by the following procedure: start with the root $ab$ and keep applying $U$ to the left and applying $V$ to the right. The set of vertices of the Cohn tree, denoted as $P$, gives a full description of periods of strongly admissible bi-infinite words.
@@
However, it is not so immediate to give examples of bi-infinite words with Markov value exactly $3$. We will give some examples later in Subsection 2.3.
@@
A bi-infinite word $\underline{\omega}$ is strongly admissible if and only if it can be written as a constant word in some alphabet $(\alpha,\beta)\in\overline{P}$, and $\underline{\omega}$ has Markov value exactly $3$ if and only if for all $n\in\mathbb{N}$, the word $\underline{\omega}$ can be written as a non-constant word for some alphabet $(\alpha_{n},\beta_{n})\in\overline{P}_{n}$.
@@
Actually Bombieri (Bom, , Theorem 15) only gave a proof that having Markov value exactly $3$ implies infinitely renormalizable. Here we give a short proof of infinitely renormalizable implies Markov value $3$, assuming Markov’s theorem.
If $\underline{\omega}$ has Markov value smaller than $3$, it will be periodic and finitely renormalizable, which contradict to the fact that $\underline{\omega}$ is infinitely renormalizable. On the other hand, if $\underline{\omega}$ has Markov value bigger than $3$, then by Lemma 2.10 there exist a bad cut of form $b\omega^{T}b|a\omega a$ in $\underline{\omega}$ or $\underline{\omega}^{T}$.
If both operators $\overline{U},\overline{V}$ appear infinitely many times in the renormalization of $\underline{\omega}$, the length of both letters $\alpha_{n},\beta_{n}$ will go to infinity, so the subword $b\omega^{T}b|a\omega a$ will be a factor of $\alpha_{N}$ or $\beta_{N}$ for some $N$ big enough. This contradicts to the fact that $\alpha_{N},\beta_{N}$ are periods of some strongly admissible words, which do not contain bad cuts.
@@
First notice that the word $\underline{\omega}$ can always be written in alphabet $\{\alpha_{n},\beta_{n}\}$ for any $n\in\mathbb{N}$ and the operators $\overline{U},\overline{V}$ are always well defined, so all the exponents of $\alpha_{n}$ or $\beta_{n}$ in $\underline{\omega}$ are $1$. In our case after the first $N_{1}$ steps, the exponents of $\beta_{n}(n\geq N_{1})$ are always $1$. We are going to show that $\underline{\omega}=\alpha_{N_{1}}^{\infty}\beta_{N_{1}}\alpha_{N_{1}}^{\infty}$ is the only possible option. If there exist some $\alpha_{N_{1}}$ with finite exponent, then under some finite step $N_{2}>N_{1}$ of renormalization the exponent will reduce to 0. But that will produce some $\beta_{N_{2}}$ such that the exponent is bigger than $1$, which means that in the step $N_{2}$ the word will have all exponents $1$ on the letter $\alpha_{N_{2}}$. Moreover, in the $N_{2}$ step of renormalization the operator will again be $\overline{U}$, which contradicts the assumption. So the exponent of $\alpha_{N_{1}}$ in $\underline{\omega}$ should be infinity, and the only possible option is the degenerate case $\alpha_{N_{1}}^{\infty}\beta_{N_{1}}\alpha_{N_{1}}^{\infty}$.
However, this again contradicts with the fact that $\underline{\omega}$ has Markov value bigger than $3$ since in this case the Markov value of $\underline{\omega}$ is going to be $3$. ∎
@@ 2.4. Properties of Renormalization
Let $\underline{\omega}$ be an infinitely renormalizable bi-infinite word ((so it has Markov value $3)$. Consider the corresponding sequence of renormalization operators $\overline{R}=\overline{R}_{1}\overline{R}_{2}\overline{R}_{3}\dots$, where $\overline{R}_{i}\in\{\overline{U},\overline{V}\}$ is the renormalization operator we applied at the $i$-th step. Then $\underline{\omega}$ has Lagrange value smaller than $3$ if and only if one of the operators $\overline{U},\overline{V}$ appears only finitely many times in $\overline{R}$. As a consequence, $\underline{\omega}$ has Lagrange value exactly $3$ if and only if both $\overline{U}$ and $\overline{V}$ appear infinitely many times in $\overline{R}$.
@@
If $\underline{\omega}$ has Lagrange value smaller than $3$, then by Markov’s Theorem it is ultimately periodic to the right with period given by the Cohn tree. So the right side of $\underline{\omega}$ will be $\alpha_{N}\beta_{N}\alpha_{N}\beta_{N}\alpha_{N}\beta_{N}\dots$ for some $\alpha_{N}\beta_{N}\in P$ under some finite step $N$ of renormalization and in the next step the right side will be $\alpha_{N+1}\alpha_{N+1}\alpha_{N+1}\dots$ or $\beta_{N+1}\beta_{N+1}\beta_{N+1}\dots$. Since $\underline{\omega}$ has Markov value $3$, the only possible option is that $\underline{\omega}$ reaches degenerate case $\alpha_{N+1}^{\infty}\beta_{N+1}\alpha_{N+1}^{\infty}$ or $\beta_{N+1}^{\infty}\alpha_{N+1}\beta_{N+1}^{\infty}$. So one of $\overline{U},\overline{V}$ can only appear at the first $N$ steps, which is finitely many times. On the other hand if one of $\overline{U},\overline{V}$ only appears for finitely many times (without loss of generality we suppose it to be $\overline{U}$), then there exist a $N_{1}$ big enough such that $\overline{U}$ only appears at the first $N_{1}$ steps. The same argument in Proposition 2.16 shows that $\underline{\omega}$ reaches the degenerate case $\beta_{N_{2}}^{\infty}\alpha_{N_{2}}\beta_{N_{2}}^{\infty}$ under some finite step $N_{2}>N_{1}$, so we have $l(\underline{\omega})=l(\beta_{N_{2}}^{\infty}\alpha_{N_{2}}\beta_{N_{2}}^{\infty})=l(\beta_{N_{2}}^{\infty})<3$. ∎
@@ 2.5. Examples of Infinitely Renormalizable Words
The periodic case corresponds to the finitely renormalizable case and the degenerate and not eventually periodic cases correspond to the infinitely renormalizable case. Moreover, $\underline{\omega}$ has Lagrange value $3$ if and only if it is infinitely renormalizable in the not eventually periodic case.
Let us give some not eventually periodic examples of bi-infinite words with Markov value exactly $3$. First we define the convergence of finite words to infinite or bi-infinite words.
@@
Given $\overline{R}=\overline{R}_{1}\overline{R}_{2}\dots$ an infinite word of renormalization operators $\overline{U},\overline{V}$ such that both the operators $\overline{U},\overline{V}$ appears infinitely many times. Let $(\alpha_{0},\beta_{0})=(a,b)$ and $(\alpha_{n+1},\beta_{n+1})=\overline{R}_{n+1}(\alpha_{n},\beta_{n})\in\overline{P}_{n+1}$ for $n\geq 0$, then sequences $\{\beta_{i}|\alpha_{i}\}_{i\in\mathbb{N}}$, $\{\alpha_{i}^{T}|\beta_{i}^{T}\}_{i\in\mathbb{N}}$ will converge to some bi-infinite words with both Lagrange value and Markov value exactly $3$.
@@ 3.1.1. Projection of infinitely renormalizable bi-infinite words
It follows from Proposition 2.26 that both of them converges, and that both the limit words $\underline{\omega}_{X_{3}(0)}=\lim_{i\to\infty}\alpha_{i}^{T}|\beta_{i}^{T}$, $\underline{\omega}_{X_{3}(\infty)}=\lim_{i\to\infty}\beta_{i}|\alpha_{i}\beta_{i}\beta_{i}$ will have Lagrange value and Markov value exactly $3$. This is because by (5) the renormalization by $\widetilde{U},\widetilde{V}$ is a particular case of renormalization by $\overline{U},\overline{V}$ and both the operators $\overline{U},\overline{V}$ will appear infinitely many times. On the other hand, clearly we have $\pi(\underline{\omega}_{X_{3}(0)})=\omega_{X_{3}(0)}$ and $\pi(\underline{\omega}_{X_{3}(\infty)})=\omega_{X_{3}(\infty)}$, which finishes the proof. ∎
@@ 3.2. Uncountably many elements in $Y_{3}(\infty)$
This lemma is a consequence of the property that $\alpha_{n}\alpha_{n}\beta_{n}\beta_{n}$ is of the form $a\hat{\theta}ab\hat{\theta}b$ with $\hat{\theta}$ palindrome. By Proposition 2.11 we know that the cut $a\hat{\theta}a|b\hat{\theta}b$ is a bad cut. So if we extend $\alpha_{n}\alpha_{n}\beta_{n}\beta_{n}$ to the right to $\alpha_{n}\omega_{X_{3}(\infty)}$, the cut $a\hat{\theta}a|b\hat{\theta}b\dots$ inside the extended word will have value bigger than $3$. On the other hand, the same cut inside the word $\omega_{X_{3}(\infty)}$ is still a bad cut. This follows from the inequality
@@
3.3. Uncountably many elements in $Y_{3}(0)$
@@
Let $\alpha_{n}^{T},\beta_{n}^{T}$ be factors of some elements in $\pi(m^{-1}(3))$ occurring at an odd position, then all the cuts inside $\alpha_{n}^{T},\beta_{n}^{T}$ are good except the cuts at the last letter $2$.
@@
First remark is that all the $\alpha_{n},\beta_{n}$ start with letter $a$, so all the $\alpha_{n}^{T},\beta_{n}^{T}$ end with letter $a$, which also means that end with letter $2$. We use induction to prove the lemma. In the case $n=0$, let $\alpha_{0}^{T}=a=22,\beta_{0}^{T}=ba=1122$ be factors of $\omega\in\pi(m^{-1}(3))$. Notice that by Proposition 2.12, the odd cuts are all good cuts, so we only need to check for the cut $1|122$, which is a good cut inside $\omega$. So the lemma holds for the case $n=0$.
@@
By induction we know that we only need to check for the cuts at the last letter $2$ of each $\alpha_{r}^{T}$ and $\beta_{r}^{T}$ (except the last $2$ of $\alpha_{r+1}^{T}$ or $\beta_{r+1}^{T}$). In general, these appear only in two situations: $(\beta_{r}^{T})^{-}a|\beta_{r}^{T}$ or $(\beta_{r}^{T})^{-}a|\alpha_{r}^{T}$. We are going to prove that those cuts are all good cuts.
  • Case $1$: $(\beta_{r}^{T})^{-}a|\beta_{r}^{T}$. Since we start with $\beta_{0}=ab$, by induction we know that there exist some $(\alpha,\beta)\in\overline{P}$ such that $\beta_{r}=\alpha\beta$. Together with (1) of Lemma 2.21 we know that we can write $\beta_{r}^{T}$ as $b\theta_{r}a$, where $\theta_{r}$ is palindrome. So the cut $(\beta_{r}^{T})^{-}a|\beta_{r}^{T}$ can be written as $b\theta_{r}a|b\theta_{r}a$, which is a good cut (Proposition 2.11).
  • Case $2$: $(\beta_{r}^{T})^{-}a|\alpha_{r}^{T}$. Since $\beta_{r}^{T}$ always ends with $\alpha_{r}^{T}$, we can write $\beta_{r}^{T}$ as $\gamma_{r}^{T}\alpha_{r}^{T}$, where $\gamma_{r}$ is some finite word. So the cut $(\beta_{r}^{T})^{-}a|\alpha_{r}^{T}$ can be written as $\gamma_{r}^{T}(\alpha_{r}^{T})^{-}a|\alpha_{r}^{T}$. Same as in the case $1$, we can write $\alpha_{r}^{T}$ as $b\hat{\theta}_{r}a$, so the cut is of the form $\gamma_{r}^{T}b\hat{\theta}_{r}a|b\hat{\theta}_{r}a$, which is also a good cut (Proposition 2.11).
@@
This follows from the Proposition 3.1 and Lemma 3.6. Suppose that there is a bad cut inside $\omega_{X_{3}(0)}$. Since $\omega_{X_{3}(0)}=\lim_{n\to\infty}\beta_{n}^{T}$, the cut must be inside (and not the cut at the last letter $2$) for some $\beta_{N}^{T}$ with $N$ big enough. But by Lemma 3.6, the cut must be a good cut inside $\omega_{X_{3}(0)}$, which is a contradiction. So $\omega_{X_{3}(0)}\in X_{3}(0)$ and by Proposition 3.1 it holds that $\omega_{X_{3}(0)}\in Y_{3}(0)$. ∎

### v3

@@ 1. Introduction
$1.$ Lagrange Spectrum and bi-infinite sequences
@@ 1.2. Statement of the main result
A classical result by Markov Markoff1879 ,Markoff1880 from 1880’s states that $\mathcal{L}$ and $\mathcal{M}$ coincide before 33 and this initial part is exactly a discrete sequence accumulating at 33, which is
@@
where $l_{1}<l_{2}<l_{3}<\dots$ is an increasing sequence of positive integers and we can show that those $y$ have Lagrange value 33.
Remark $1.$
@@
It is relatively easy to give concrete examples of sequences in $m^{-1}(3)$, but these “easy” examples are periodic at both sides and have Lagrange value less than 33. Sequences in $m^{-1}(3)\cap l^{-1}(3)$ are not so direct to describe, nevertheless with the help of Bombieri’s characterization we have a criteria (see Lemma 2.22) to produce quite a lot such elements.
@@
A detailed sketch of proof is provided at the beginning of Section 33. We first construct uncountably many elements in the set $Y_{3}(\infty)$ and $Y_{3}(0)$. Then by gluing some carefully chosen finite words at the beginning of elements in $Y_{3}(0)$, we can map $Y_{3}(0)$ injectively into $Y_{3}(n)$ for $0<n<\infty$.
@@ 2.1. Markov’s Theorem
A well known uniqueness conjecture by Frobenius states that all the Markov numbers have multiplicity of 11 and the conjecture still remains open.
Markov’s theorem gives a complete description of the spectrum $\mathcal{L}$ and $\mathcal{M}$ at the part before 33. Here we state a version in Bom .
@@
$2.$ Structure of Admissible Bi-infinite Words
@@
If there is one letter bigger than 22 appearing in a bi-infinite word, then the Markov value will be strictly bigger than 33, so the only possible letters that could appear in an admissible bi-infinite word are 11 and 22. Moreover, it was shown in (Bom, , Lemma 9) that the number of consecutive 11s or 22s in an admissible bi-infinite word is always even or infinite, so an admissible bi-infinite word can be written as a bi-infinite word of alphabet $\{a=22,b=11\}$. In (Bom, , Lemma 11), admissible bi-infinite words were classified into four possible types:
@@
We only consider positive words, that is, words with positive exponents. If the word is no longer positive after applying $u$ or $v$, then we say that the substitution $u$ or $v$ is not well defined. A remark is that if $u$ or $v$ is well defined for some bi-infinite word $\underline{\omega}$, then all the exponents of $a$ or $b$ in the word $\underline{\omega}$ should be 11.
@@
We will focus on cuts with value bigger than 33 or not bigger than 33.
@@
A cut of an infinite or bi-infinite word is bad if the cut has value strictly bigger than 33, otherwise the cut is good.
@@
For every $n\in\mathbb{N}\cup\{\infty\}$, the elements in $X_{3}(n)$ are infinite words with Lagrange value 33 and with exactly $n$ bad cuts. As was stated in Section 1, our main idea is to take projections of elements in $m^{-1}(3)\cap l^{-1}(3)$, and we have to control the number of bad cuts inside the projections.
@@
An important observation is that good cuts for a bi-infinite word $\underline{\omega}$ can change to bad for the infinite word $\pi(\underline{\omega})$. For example, the bi-infinite word $(abb)^{\infty}$ has only good cuts (since it has Markov value less than 33) but the projection $abbabb\dots$ has bad cuts, because
@@ 2.3. Renormalization
A non-constant bi-infinite word $\underline{\omega}$ is strongly admissible if and only if the renormalization algorithm stops at some finite step and reaches $\underline{\omega}_{0}$, and a non-constant bi-infinite word has Markov value exactly 33 if and only if the renormalization algorithm never stops.
@@
The Cohn tree $T$ is an infinite binary tree generated by the following procedure: start with the root abab and keep applying $U$ to the left and applying $V$ to the right. The set of vertices of the Cohn tree, denoted as $P$, gives a full description of periods of strongly admissible bi-infinite words.
@@
However, it is not so immediate to give examples of bi-infinite words with Markov value exactly 33. We will give some examples later in Subsection 2.3.
@@
A bi-infinite word $\underline{\omega}$ is strongly admissible if and only if it can be written as a constant word in some alphabet $(\alpha,\beta)\in\overline{P}$, and $\underline{\omega}$ has Markov value exactly 33 if and only if for all $n\in\mathbb{N}$, the word $\underline{\omega}$ can be written as a non-constant word for some alphabet $(\alpha_{n},\beta_{n})\in\overline{P}_{n}$.
@@
Actually Bombieri (Bom, , Theorem 15) only gave a proof that having Markov value exactly 33 implies infinitely renormalizable. Here we give a short proof of infinitely renormalizable implies Markov value 33, assuming Markov’s theorem.
If $\underline{\omega}$ has Markov value smaller than 33, it will be periodic and finitely renormalizable, which contradict to the fact that $\underline{\omega}$ is infinitely renormalizable. On the other hand, if $\underline{\omega}$ has Markov value bigger than 33, then by Lemma 2.10 there exist a bad cut of form $b\omega^{T}b|a\omega a$ in $\underline{\omega}$ or $\underline{\omega}^{T}$.
If both operators $\overline{U},\overline{V}$ appear infinitely many times in the renormalization of $\underline{\omega}$, the length of both letters $\alpha_{n},\beta_{n}$ will go to infinity, so the subword $b\omega^{T}b|a\omega a$ will be a factor of $\alpha_{N}$ or $\beta_{N}$ for some NN big enough. This contradicts to the fact that $\alpha_{N},\beta_{N}$ are periods of some strongly admissible words, which do not contain bad cuts.
@@
First notice that the word $\underline{\omega}$ can always be written in alphabet $\{\alpha_{n},\beta_{n}\}$ for any $n\in\mathbb{N}$ and the operators $\overline{U},\overline{V}$ are always well defined, so all the exponents of $\alpha_{n}$ or $\beta_{n}$ in $\underline{\omega}$ are 11. In our case after the first $N_{1}$ steps, the exponents of $\beta_{n}(n\geq N_{1})$ are always 11. We are going to show that $\underline{\omega}=\alpha_{N_{1}}^{\infty}\beta_{N_{1}}\alpha_{N_{1}}^{\infty}$ is the only possible option. If there exist some $\alpha_{N_{1}}$ with finite exponent, then under some finite step $N_{2}>N_{1}$ of renormalization the exponent will reduce to 0. But that will produce some $\beta_{N_{2}}$ such that the exponent is bigger than 11, which means that in the step $N_{2}$ the word will have all exponents 11 on the letter $\alpha_{N_{2}}$. Moreover, in the $N_{2}$ step of renormalization the operator will again be $\overline{U}$, which contradicts the assumption. So the exponent of $\alpha_{N_{1}}$ in $\underline{\omega}$ should be infinity, and the only possible option is the degenerate case $\alpha_{N_{1}}^{\infty}\beta_{N_{1}}\alpha_{N_{1}}^{\infty}$.
However, this again contradicts with the fact that $\underline{\omega}$ has Markov value bigger than 33 since in this case the Markov value of $\underline{\omega}$ is going to be 33. ∎
@@ 2.4. Properties of Renormalization
Let $\underline{\omega}$ be an infinitely renormalizable bi-infinite word ((so it has Markov value $3)$. Consider the corresponding sequence of renormalization operators $\overline{R}=\overline{R}_{1}\overline{R}_{2}\overline{R}_{3}\dots$, where $\overline{R}_{i}\in\{\overline{U},\overline{V}\}$ is the renormalization operator we applied at the $i$-th step. Then $\underline{\omega}$ has Lagrange value smaller than 33 if and only if one of the operators $\overline{U},\overline{V}$ appears only finitely many times in $\overline{R}$. As a consequence, $\underline{\omega}$ has Lagrange value exactly 33 if and only if both $\overline{U}$ and $\overline{V}$ appear infinitely many times in $\overline{R}$.
@@
If $\underline{\omega}$ has Lagrange value smaller than 33, then by Markov’s Theorem it is ultimately periodic to the right with period given by the Cohn tree. So the right side of $\underline{\omega}$ will be $\alpha_{N}\beta_{N}\alpha_{N}\beta_{N}\alpha_{N}\beta_{N}\dots$ for some $\alpha_{N}\beta_{N}\in P$ under some finite step NN of renormalization and in the next step the right side will be $\alpha_{N+1}\alpha_{N+1}\alpha_{N+1}\dots$ or $\beta_{N+1}\beta_{N+1}\beta_{N+1}\dots$. Since $\underline{\omega}$ has Markov value 33, the only possible option is that $\underline{\omega}$ reaches degenerate case $\alpha_{N+1}^{\infty}\beta_{N+1}\alpha_{N+1}^{\infty}$ or $\beta_{N+1}^{\infty}\alpha_{N+1}\beta_{N+1}^{\infty}$. So one of $\overline{U},\overline{V}$ can only appear at the first NN steps, which is finitely many times. On the other hand if one of $\overline{U},\overline{V}$ only appears for finitely many times (without loss of generality we suppose it to be $\overline{U}$), then there exist a $N_{1}$ big enough such that $\overline{U}$ only appears at the first $N_{1}$ steps. The same argument in Proposition 2.16 shows that $\underline{\omega}$ reaches the degenerate case $\beta_{N_{2}}^{\infty}\alpha_{N_{2}}\beta_{N_{2}}^{\infty}$ under some finite step $N_{2}>N_{1}$, so we have $l(\underline{\omega})=l(\beta_{N_{2}}^{\infty}\alpha_{N_{2}}\beta_{N_{2}}^{\infty})=l(\beta_{N_{2}}^{\infty})<3$. ∎
@@ 2.5. Examples of Infinitely Renormalizable Words
The periodic case corresponds to the finitely renormalizable case and the degenerate and not eventually periodic cases correspond to the infinitely renormalizable case. Moreover, $\underline{\omega}$ has Lagrange value 33 if and only if it is infinitely renormalizable in the not eventually periodic case.
Let us give some not eventually periodic examples of bi-infinite words with Markov value exactly 33. First we define the convergence of finite words to infinite or bi-infinite words.
@@
Given $\overline{R}=\overline{R}_{1}\overline{R}_{2}\dots$ an infinite word of renormalization operators $\overline{U},\overline{V}$ such that both the operators $\overline{U},\overline{V}$ appears infinitely many times. Let $(\alpha_{0},\beta_{0})=(a,b)$ and $(\alpha_{n+1},\beta_{n+1})=\overline{R}_{n+1}(\alpha_{n},\beta_{n})\in\overline{P}_{n+1}$ for $n\geq 0$, then sequences $\{\beta_{i}|\alpha_{i}\}_{i\in\mathbb{N}}$, $\{\alpha_{i}^{T}|\beta_{i}^{T}\}_{i\in\mathbb{N}}$ will converge to some bi-infinite words with both Lagrange value and Markov value exactly 33.
@@ 3.1.1. Projection of infinitely renormalizable bi-infinite words
It follows from Proposition 2.26 that both of them converges, and that both the limit words $\underline{\omega}_{X_{3}(0)}=\lim_{i\to\infty}\alpha_{i}^{T}|\beta_{i}^{T}$, $\underline{\omega}_{X_{3}(\infty)}=\lim_{i\to\infty}\beta_{i}|\alpha_{i}\beta_{i}\beta_{i}$ will have Lagrange value and Markov value exactly 33. This is because by (5) the renormalization by $\widetilde{U},\widetilde{V}$ is a particular case of renormalization by $\overline{U},\overline{V}$ and both the operators $\overline{U},\overline{V}$ will appear infinitely many times. On the other hand, clearly we have $\pi(\underline{\omega}_{X_{3}(0)})=\omega_{X_{3}(0)}$ and $\pi(\underline{\omega}_{X_{3}(\infty)})=\omega_{X_{3}(\infty)}$, which finishes the proof. ∎
@@ 3.2. Uncountably many elements in $Y_{3}(\infty)$
This lemma is a consequence of the property that $\alpha_{n}\alpha_{n}\beta_{n}\beta_{n}$ is of the form $a\hat{\theta}ab\hat{\theta}b$ with $\hat{\theta}$ palindrome. By Proposition 2.11 we know that the cut $a\hat{\theta}a|b\hat{\theta}b$ is a bad cut. So if we extend $\alpha_{n}\alpha_{n}\beta_{n}\beta_{n}$ to the right to $\alpha_{n}\omega_{X_{3}(\infty)}$, the cut $a\hat{\theta}a|b\hat{\theta}b\dots$ inside the extended word will have value bigger than 33. On the other hand, the same cut inside the word $\omega_{X_{3}(\infty)}$ is still a bad cut. This follows from the inequality
@@
$3.$ Uncountably many elements in $Y_{3}(0)$
@@
Let $\alpha_{n}^{T},\beta_{n}^{T}$ be factors of some elements in $\pi(m^{-1}(3))$ occurring at an odd position, then all the cuts inside $\alpha_{n}^{T},\beta_{n}^{T}$ are good except the cuts at the last letter 22.
@@
First remark is that all the $\alpha_{n},\beta_{n}$ start with letter $a$, so all the $\alpha_{n}^{T},\beta_{n}^{T}$ end with letter $a$, which also means that end with letter 22. We use induction to prove the lemma. In the case $n=0$, let $\alpha_{0}^{T}=a=22,\beta_{0}^{T}=ba=1122$ be factors of $\omega\in\pi(m^{-1}(3))$. Notice that by Proposition 2.12, the odd cuts are all good cuts, so we only need to check for the cut $1|122$, which is a good cut inside $\omega$. So the lemma holds for the case $n=0$.
@@
By induction we know that we only need to check for the cuts at the last letter 22 of each $\alpha_{r}^{T}$ and $\beta_{r}^{T}$ (except the last 22 of $\alpha_{r+1}^{T}$ or $\beta_{r+1}^{T}$). In general, these appear only in two situations: $(\beta_{r}^{T})^{-}a|\beta_{r}^{T}$ or $(\beta_{r}^{T})^{-}a|\alpha_{r}^{T}$. We are going to prove that those cuts are all good cuts.
  • Case 11: $(\beta_{r}^{T})^{-}a|\beta_{r}^{T}$. Since we start with $\beta_{0}=ab$, by induction we know that there exist some $(\alpha,\beta)\in\overline{P}$ such that $\beta_{r}=\alpha\beta$. Together with (1) of Lemma 2.21 we know that we can write $\beta_{r}^{T}$ as $b\theta_{r}a$, where $\theta_{r}$ is palindrome. So the cut $(\beta_{r}^{T})^{-}a|\beta_{r}^{T}$ can be written as $b\theta_{r}a|b\theta_{r}a$, which is a good cut (Proposition 2.11).
  • Case 22: $(\beta_{r}^{T})^{-}a|\alpha_{r}^{T}$. Since $\beta_{r}^{T}$ always ends with $\alpha_{r}^{T}$, we can write $\beta_{r}^{T}$ as $\gamma_{r}^{T}\alpha_{r}^{T}$, where $\gamma_{r}$ is some finite word. So the cut $(\beta_{r}^{T})^{-}a|\alpha_{r}^{T}$ can be written as $\gamma_{r}^{T}(\alpha_{r}^{T})^{-}a|\alpha_{r}^{T}$. Same as in the case 11, we can write $\alpha_{r}^{T}$ as $b\hat{\theta}_{r}a$, so the cut is of the form $\gamma_{r}^{T}b\hat{\theta}_{r}a|b\hat{\theta}_{r}a$, which is also a good cut (Proposition 2.11).
@@
This follows from the Proposition 3.1 and Lemma 3.6. Suppose that there is a bad cut inside $\omega_{X_{3}(0)}$. Since $\omega_{X_{3}(0)}=\lim_{n\to\infty}\beta_{n}^{T}$, the cut must be inside (and not the cut at the last letter 22) for some $\beta_{N}^{T}$ with NN big enough. But by Lemma 3.6, the cut must be a good cut inside $\omega_{X_{3}(0)}$, which is a contradiction. So $\omega_{X_{3}(0)}\in X_{3}(0)$ and by Proposition 3.1 it holds that $\omega_{X_{3}(0)}\in Y_{3}(0)$. ∎